# When Bias Pretends to Be Truth: How Spurious Correlations Undermine Hallucination Detection in LLMs

### v1

@@ Appendix B Proofs
The kernel function $k(x,x^{\prime})$ is positive definite for any $x,x^{\prime}\in\mathcal{X}$. The objective function $f\in\mathcal{H}_{k}$ lives in the reproducing kernel Hilbert space (RKHS) induced by $k$. The RKHS endowed with the inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}_{k}}$ is defined as
$\mathcal{H}_{k}\coloneqq\Biggl\{f=\sum_{i=1}^{\infty}c_{i}k(\cdot,x_{i})\;:$ $(c_{1},c_{2},\dots)\subset\mathbb{R},\quad(x_{1},x_{2},\dots)\subset\mathcal{X},\quad\text{ such that}$
$\|f\|^{2}_{\mathcal{H}_{k}}\coloneqq\lim_{n\to\infty}\left\|\sum_{i=1}^{n}c_{i}k(\cdot,x_{i})\right\|^{2}_{\mathcal{H}_{k}}=\sum_{i,j=1}^{\infty}c_{i}c_{j}k(x_{i},x_{j})<\infty\Biggr\},$
and for any $f,g\in\mathcal{H}_{k}$ with $f=\sum_{i=1}^{\infty}c_{i}k(\cdot,x_{i})$ and $g=\sum_{j=1}^{\infty}c_{j}^{\prime}k(\cdot,x_{j}^{\prime})$,
$\langle f,g\rangle_{\mathcal{H}_{k}}\coloneqq\sum_{i,j=1}^{\infty}c_{i}c_{j}^{\prime}k(x_{i},x_{j}^{\prime}).$
@@
The target function $f^{*}$ lies in the RKHS $\mathcal{H}_{k}(\mathbb{R}^{d})$, with $\|f^{*}\|_{\mathcal{H}_{k}}\leq B$ for some constant $B>0$.
@@
For the case of interpolation, the kernel matrix $k(X_{N},X_{N})$ is invertible.
Note that dot-product kernels on $\mathbb{S}^{d}$ are all radius basis function kernels, hence translation-invariant. So that the above assumptions are also satisfied by neural kernels such as NNGP and NTK. Assumption 3 avoids the Gibbs phenomenon at the boundary between the correlation and noisy regions, ensuring that the target function can be well approximated by functions in the RKHS. Assumption 4 guarantees the uniqueness of the kernel interpolation solution.
@@ B.1 Kernel Ridge Regression with Fixed Bandwidth
Under Assumptions 1-3, there exist some constants $C_{0},C_{1},C_{2},C_{3},N_{0}>0$ such that, for any $\delta\in(0,1)$ and $N\geq N_{0}$, define the uniform upper confidence bound as
$U_{N}^{\delta}\coloneqq C_{0}C_{2}^{1/2}\left(\frac{\ln(N/2)}{C_{3}N}\right)^{\frac{2s-d-1}{2d}}\sqrt{\ln(1+\lambda N)\ln(e+2C_{1}/\delta)}.$
@@
Therefore, for any threshold $\rho\in(0,0.98)$, if $N$ is sufficiently large, then
$\mathbb{P}\left(\left\{|f_{N}(x)|\geq\rho,\kern 5.0pt\forall x\in\mathcal{C}\right\}\cap\left\{|f_{N}(x^{\prime})|<\rho,\kern 5.0pt\forall x^{\prime}\in\mathcal{N}\right\}\right)\geq 1-\delta,$
which indicates that any hallucination detection criterion with a fixed $\rho$ fails in both the correlation and noisy regions.
@@
Under Assumptions 1-3, for any $N\geq 1$, $\delta\in(0,1)$, $\left\{x_{1},\dots,x_{N}\right\}\subset\mathbb{S}^{d}$, and independent sub-Gaussian random variables $\left\{\varepsilon_{1},\dots,\epsilon_{N}\right\}$ with mean zero and variance proxy $\varsigma^{2}$, there exist constants $C_{0},C_{1}>0$ only depending on $k$, $d$ and $\varsigma$ such that
@@
The original statement in wang2023regret holds when the domain is assumed to be compact and convex. The convexity assumption can be removed as follows: First, a classical approach is to extend the domain to a compact set with Lipschitz boundary and satisfying the interior cone condition (wendland2004scattered). Second, by the Sobolev Extension Theorem (mclean2000strongly), any function in $W_{2}^{s}(\mathbb{S}^{d})$ can be extended to a function in $W_{2}^{s+1/2}(B_{d+1})$, where $B_{d+1}$ is the unit ball containing $\mathbb{S}^{d}$. The extension operator is linear and bounded, so the norm equivalence in Assumption 2 still holds up to a constant. Therefore, the proof of Theorem 1 in wang2023regret remains valid. ∎
@@
Lemma 4 (Theorem 5 in wu1993local).
Under Assumption 2, there exist constants $C_{2},h_{0}>0$ such that, for an arbitrary dataset $X_{N}=\left\{x_{1},\dots,x_{N}\right\}$ satisfying $h_{N}\leq h_{0}$,
$\sigma_{N}^{2}(x)\leq C_{2}h_{N}^{2s-d-1}.$
@@
Suppose $X_{N}=\left\{x_{1},\dots,x_{N}\right\}$ are independently uniformly sampled from $\mathbb{S}^{d}$. There exist a constant $C_{3}$ only depends on $d$ such that, for any $\delta\in(0,1)$ and $N\geq 3$, with probability at least $1-\delta$,
$h_{N}\leq\left(\frac{\ln(N\delta)}{C_{3}N}\right)^{1/d}.$
@@
$\leq\mathrm{card}(\mathscr{E}_{t/2})\left(1-\frac{\omega_{d}(t/4)^{d}}{\mathscr{H}_{d}(\mathbb{S}^{d})}\right)^{N},$
where $\mathrm{card}(\mathscr{E}_{\cdot})$ is the packing number satisfying
$\mathscr{H}_{d}(\mathbb{S}^{d})\geq\sum_{x\in\mathscr{E}_{t/2}}\mathscr{H}_{d}(B(x,t/4)\cap\mathbb{S}^{d})\geq\mathrm{card}(\mathscr{E}_{t/2})\omega_{d}(t/4)^{d}.$
@@
$=(Ct^{d})^{-1}\exp(-Ct^{d}N).$
Where $C\coloneqq\omega_{d}/(4^{d}\mathscr{H}_{d}(\mathbb{S}^{d}))$ is a positive constant only depending on $d$.
Let $\delta=(Ct^{d})^{-1}\exp(-Ct^{d}N)$, then $t=\left(\mathrm{W}(N/\delta)/CN\right)^{1/d}$, where $\mathrm{W}(\cdot)$ is the Lambert W function. Note that $\mathrm{W}(x)<\ln(x)$ when $x>e$, so if $N>\delta e$, then with probability at least $1-\delta$,
$h_{N}\leq\left(\frac{\ln(N\delta)}{CN}\right)^{1/d},$
@@
$h_{N}^{d}\leq\frac{\ln(N\delta/2)}{C_{3}N}\leq\frac{\ln(N/2)}{C_{3}N}.$
To satisfy the condition $h_{N}\leq h_{0}$ in Lemma 4, by the monotonicity of $\ln(x)/x$, it suffices to take $N\geq N_{0}\coloneqq\max\left\{2\ln(1/(2C_{3}h_{0}^{d}))/C_{3}h_{0}^{d},3\right\}$. Conditioned on the above $X_{N}$, by Lemma 3 and Lemma 4, with probability at least $1-\delta/2$, the following holds for all $x\in\mathbb{S}^{d}$,
@@
$\leq C_{0}C_{2}^{1/2}\left(\frac{\ln(N/2)}{C_{3}N}\right)^{\frac{2s-d-1}{2d}}\sqrt{\ln(1+\lambda N)\ln(e+2C_{1}/\delta)}\coloneqq U_{N}^{\delta}.$
@@ B.2 Kernel Ridge Regression with Decaying Bandwidth
Under the same conditions as Lemma 5, there exist constants $C_{1},C_{2}$ only depends on $d$ such that,
@@
If we set the bandwidth of KRR sufficiently small, the model is forced to memorize all data points. However, such a predictor is inconsistent (i.e., excess risk is bounded away from zero). The following Theorem 7 gives an intuitive explanation for this phenomenon.
@@
Suppose Assumptions 1-3 hold. Further, assume the kernel function has compact support, define as $k_{\ell_{N}}(x,x^{\prime})\coloneqq\Psi((x-x^{\prime})/\ell_{N})$, where $\Psi$ is supported on $B(0,1)$ (e.g., Wendland kernel (wendland1995piecewise)) and $\ell_{N}>0$ is the bandwidth. Let $\ell_{N}=o(N^{1/d})$, there exists a constant $\rho\in(0,1)$ depending on $\lambda$ such that $\lim_{\lambda\to 0}\rho=1$ and
$\lim_{N\to\infty}\mathbb{P}\left(|f_{N}(x_{i})|\geq\rho\right)=1,\quad\forall i=1,\dots,N.$
However,
$\lim_{N\to\infty}\mathbb{P}\left(f_{N}(x)\neq 0\right)=0,$
@@
By Lemma 6, for sufficiently large $N$, we have $\ell_{N}<q_{N}$ holds with probability $1$. Therefore, the kernel matrix $k_{\ell_{N}}(X_{N},X_{N})$ is diagonal dominant with diagonal entries being $\Psi(0)$ and off-diagonal entries being zero. So that
$f_{N}(x_{i})=k_{\ell_{N}}(x_{i},X_{N})(k_{\ell_{N}}(X_{N},X_{N})+N\lambda I_{N})^{-1}Y_{N}=\frac{\Psi(0)}{\Psi(0)+N\lambda}y_{i}.$
Let $\rho=|\Psi(0)|/(|\Psi(0)|+\lambda)$, then $|f_{N}(x_{i})|\geq\rho$ for all $i=1,\dots,N$.
On the other hand, the result is direct by the compact support of the kernel and the Ahlfors-David regularity,
$\mathbb{P}\left(f_{N}(x)\neq 0\right)\leq N\mu(B(x,\ell_{N}))\to 0,\quad\text{ as }N\to\infty.$
@@ B.3 Kernel Ridgeless Regression with Benign Overfitting
Suppose Assumptions 1-4 hold. Further, assume either
@@
  • the kernel function is define as $k_{c_{N},\gamma_{N}}(x,x^{\prime})\coloneqq\tilde{k}(x,x^{\prime})+c_{N}\check{k}_{\gamma_{N}}(x,x^{\prime})$, where $\tilde{k}$ is a universal kernel, $\check{k}_{\gamma_{N}}$ is the Laplace kernel with bandwidth $\gamma_{N}>0$, $c_{N}\to 0$, $Nc_{N}^{4}\to\infty$, and $\gamma_{N}\leq N^{-3/d}(7\ln N)^{-1}$.
Then for any $\delta\in(0,1)$, there exist constants $C_{0},N_{0},\alpha>0$ such that for any $N\geq N_{0}$, if we define the uniform upper confidence bound as $U_{N}^{\delta}\coloneqq C_{0}N^{-\alpha}$, the following holds
@@
Therefore, for any threshold $\rho\in(0,0.98)$,
$\liminf_{N\to\infty}\mathbb{P}\left(\mathbb{E}_{D_{N}}\left[|f_{N}(x)|\geq\rho\right]\right)\geq 1-\delta,\quad\text{ for all }x\in\mathcal{C},$
which indicates that any hallucination detection criterion with a fixed $\rho$ fails in the correlation regions.
@@
$\mathcal{E}_{N}=O_{\mathbb{P}}\left(N^{-\alpha}\right)\quad\text{ as }N,d\to\infty.$
@@
The proof is completed by combining the above result with Proposition 9 or Proposition 10. ∎
Appendix C Neural Kernels and Kernel Regime
For an over-parametrized neural network of any architecture (e.g., multi-layer perceptron (MLP), residual network (ResNet), convolutional neural network (CNN), Transformer) under standard initialization (Xavier or He initialization) or neural tangent parametrization (jacot2018neural), the stochastic gradient descent (SGD) training dynamics of its output $f:\mathbb{R}^{d}\to\mathbb{R}$ can be tracked by the kernel gradient descent

### v2

@@ Appendix B Proofs
The kernel function $k(x,x^{\prime})$ is positive definite for any $x,x^{\prime}\in\mathcal{X}$. The objective function $f\in\mathcal{H}_{k}(\mathcal{X})$ lives in the reproducing kernel Hilbert space (RKHS) induced by $k$. The RKHS endowed with the inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}_{k}(\mathcal{X})}$ is defined as
$\mathcal{H}_{k}(\mathcal{X})\coloneqq\Biggl\{f=\sum_{i=1}^{\infty}$ $c_{i}k(\cdot,x_{i})\;:\;(c_{1},c_{2},\dots)\subset\mathbb{R},\quad(x_{1},x_{2},\dots)\subset\mathcal{X},\quad\text{ such that}$
$\|f\|^{2}_{\mathcal{H}_{k}(\mathcal{X})}\coloneqq\lim_{n\to\infty}\left\|\sum_{i=1}^{n}c_{i}k(\cdot,x_{i})\right\|^{2}_{\mathcal{H}_{k}(\mathcal{X})}=\sum_{i,j=1}^{\infty}c_{i}c_{j}k(x_{i},x_{j})<\infty\Biggr\},$
and for any $f,g\in\mathcal{H}_{k}(\mathcal{X})$ with $f=\sum_{i=1}^{\infty}c_{i}k(\cdot,x_{i})$ and $g=\sum_{j=1}^{\infty}c_{j}^{\prime}k(\cdot,x_{j}^{\prime})$,
$\langle f,g\rangle_{\mathcal{H}_{k}(\mathcal{X})}\coloneqq\sum_{i,j=1}^{\infty}c_{i}c_{j}^{\prime}k(x_{i},x_{j}^{\prime}).$
@@
The target function $f^{*}$ lies in the RKHS $\mathcal{H}_{k}(\mathbb{S}^{d})$, with $\|f^{*}\|_{\mathcal{H}_{k}(\mathbb{S}^{d})}\leq B$ for some constant $B>0$.
@@
The kernel matrix $k(X_{N},X_{N})$ is invertible.
Assumptions 1 and 2 are standard in the analysis of kernel ridge regression. Since dot-product kernels on $\mathbb{S}^{d}$ are radial basis functions and thus translation-invariant, these assumptions also hold for neural kernels such as NNGP and NTK (see Appendix C for further details). Assumption 3 avoids the Gibbs phenomenon at the boundary between the correlation and noisy regions, ensuring that the target function can be well approximated by functions in the RKHS. Assumption 4 guarantees the distinctness of all data points in $X_{N}$ and ensures the uniqueness of the kernel interpolation solution.
@@ B.1 Kernel Ridge Regression with Fixed Bandwidth
Under Assumptions 1-3, there exist constants $C_{0},C_{1},C_{2},C_{3}>0$ such that for any $\delta\in(0,1)$ and $N\geq N_{0}$ with $N_{0}=O(\ln(1/\delta))$, define the uniform upper confidence bound as
$U_{N}^{\delta}\coloneqq C_{0}C_{2}^{1/2}\left(\frac{\ln(2N/\delta)}{C_{3}N}\right)^{\frac{2s-d-1}{2d}}\sqrt{\ln(1+\lambda N)\ln(e+2C_{1}/\delta)}.$
@@
Therefore, for any threshold $\tau\in(0,0.98)$, if $N$ is sufficiently large, then
$\mathbb{P}\left(\left\{|f_{N}(x)|\geq\tau,\kern 5.0pt\forall x\in\mathcal{C}\right\}\cap\left\{|f_{N}(x^{\prime})|<\tau,\kern 5.0pt\forall x^{\prime}\in\mathcal{N}\right\}\right)\geq 1-\delta,$
which indicates that any hallucination detection criterion with a fixed $\tau$ fails in both the correlation and noisy regions.
@@
Under Assumptions 1-3, for any $N\geq 1$, $\delta\in(0,1)$, $\left\{x_{1},\dots,x_{N}\right\}\subset\mathbb{S}^{d}$, and independent sub-Gaussian random variables $\left\{\varepsilon_{1},\dots,\epsilon_{N}\right\}$ with mean zero and variance proxy $\varsigma^{2}$, there exist constants $C_{0},C_{1}>0$ only depending on $k$, $d$, $B$, $\nu$ and $\varsigma^{2}$ such that
@@
The original statement in wang2023regret holds when the domain is assumed to be compact and convex. The convexity assumption can be removed as follows: First, a classical approach is to extend the domain to a compact set with Lipschitz boundary and satisfying the interior cone condition (wendland2004scattered). Second, by the Sobolev extension theorem (mclean2000strongly), any function in $W_{2}^{s}(\mathbb{S}^{d})$ can be extended to a function in $W_{2}^{s+1/2}(\bar{B}_{d+1})$, where $\bar{B}_{d+1}$ is the unit closed ball such that $\mathbb{S}^{d}=\partial B_{d+1}\subset\bar{B}_{d+1}$. The extension operator is linear and bounded, so the norm equivalence in Assumption 2 still holds up to a constant. Therefore, the proof of Theorem 1 in wang2023regret remains valid. ∎
@@
Lemma 4 (Theorem 5 in wu1993local; Theorem 5.4 in kanagawa2018gaussian).
Under Assumption 2, there exist constants $C_{2},h_{0}>0$ such that, for an arbitrary dataset $X_{N}=\left\{x_{1},\dots,x_{N}\right\}\subset\mathbb{S}^{d}$ satisfying $h_{N}\leq h_{0}$,
$\sigma_{N}^{2}(x)\leq C_{2}h_{N}^{2s-d-1},\quad\text{ for all }x\in\mathbb{S}^{d}.$
@@
Suppose $X_{N}=\left\{x_{1},\dots,x_{N}\right\}$ are independently uniformly sampled from $\mathbb{S}^{d}$. There exist a constant $C_{3}$ only depending on $d$ such that, for any $\delta\in(0,1)$ and $N\geq 3$, with probability at least $1-\delta$,
$h_{N}\leq\left(\frac{\ln(N/\delta)}{C_{3}N}\right)^{1/d}.$
@@
$\leq\#(\mathscr{E}_{t/2})\left(1-\frac{\omega_{d}(t/4)^{d}}{\mathscr{H}_{d}(\mathbb{S}^{d})}\right)^{N},$
where $(\mathscr{E}_{t/2})$ is the $t/2$-packing number of $\mathbb{S}^{d}$ satisfying
$\mathscr{H}_{d}(\mathbb{S}^{d})\geq\sum_{x\in\mathscr{E}_{t/2}}\mathscr{H}_{d}(B(x,t/4)\cap\mathbb{S}^{d})\geq\#(\mathscr{E}_{t/2})\omega_{d}(t/4)^{d}.$
@@
$=(C_{3}t^{d})^{-1}\exp(-C_{3}t^{d}N).$
Where $C_{3}\coloneqq\omega_{d}/(4^{d}\mathscr{H}_{d}(\mathbb{S}^{d}))$ is a positive constant only depending on $d$.
Let $\delta=(C_{3}t^{d})^{-1}\exp(-C_{3}t^{d}N)$, then $t=\left(\mathrm{W}(N/\delta)/(C_{3}N)\right)^{1/d}$, where $\mathrm{W}(\cdot)$ is the Lambert W function. Note that $\mathrm{W}(x)<\ln(x)$ when $x>e$, so if $N>\delta e$, then with probability at least $1-\delta$,
$h_{N}\leq\left(\frac{\ln(N/\delta)}{C_{3}N}\right)^{1/d},$
@@
$h_{N}^{d}\leq\frac{\ln(2N/\delta)}{C_{3}N}.$
To satisfy the condition $h_{N}\leq h_{0}$ in Lemma 4, by the monotonicity of $\ln(x)/x$ at $e,\infty)$, it suffices to set $N\geq N_{0}\coloneqq\max\left\{2\ln(2/(C_{3}h_{0}^{d}\delta))/(C_{3}h_{0}^{d}),3\right\}$. Conditioned on the above $X_{N}$, by Lemma 3 and Lemma 4, with probability at least $1-\delta/2$, the following holds for all $x\in\mathbb{S}^{d}$,
@@
$\leq C_{0}C_{2}^{1/2}\left(\frac{\ln(2N/\delta)}{C_{3}N}\right)^{\frac{2s-d-1}{2d}}\sqrt{\ln(1+\lambda N)\ln(e+2C_{1}/\delta)}\coloneqq U_{N}^{\delta}.$
@@ B.2 Kernel Ridge Regression with Decaying Bandwidth
Under the same conditions as Lemma 5, there exist constants $C_{1},C_{2}$ only depending on $d$ such that,
@@
If we set the bandwidth of KRR sufficiently small, the model learns nothing but memorizes all data points, which results in the excess risk being bounded away from zero. The following Theorem 7 provides an intuitive explanation for this phenomenon.
@@
Under Assumption 4, and suppose the kernel function has compact support, define as $k_{\ell_{N}}(x,x^{\prime})\coloneqq\Psi((x-x^{\prime})/\ell_{N})$, where $\ell_{N}>0$ is the bandwidth, $\Psi$ is supported on $B(0,1)$ and $\Psi(0)>0$. Let $\ell_{N}=o(N^{-1/d})$, then
$\lim_{N\to\infty}\mathbb{P}\left(|f_{N}(x_{i})|\geq\frac{|\Psi(0)|}{|\Psi(0)|+\lambda N},\quad\text{ for all }i=1,\dots,N\right)=1,$
which implies that the model is able to memorize the data with a weak regularizer $\lambda=O(N^{-1})$. However,
$\lim_{N\to\infty}\mathbb{P}\left(f_{N}(x)\neq 0\right)=0,\quad\text{ for all }x\in\mathcal{X}\setminus X_{N},$
@@
By Lemma 6, for sufficiently large $N$, we have $\ell_{N}<q_{N}$ holds with probability $1$. Therefore, the kernel matrix $k_{\ell_{N}}(X_{N},X_{N})$ is diagonally dominant with diagonal entries being $\Psi(0)$ and off-diagonal entries being zero. Hence, the following holds almost surely for all $i=1,\dots,N$:
$f_{N}(x_{i})=k_{\ell_{N}}(x_{i},X_{N})(k_{\ell_{N}}(X_{N},X_{N})+\lambda NI_{N})^{-1}Y_{N}=\frac{\Psi(0)}{\Psi(0)+\lambda N}y_{i}.$
For the second part, the result follows directly by applying the compact support of the kernel and the Ahlfors-David regularity of $\mathbb{S}^{d}$,
$\mathbb{P}\left(f_{N}(x)\neq 0\right)\leq\mathbb{P}\left(\bigcup_{i=1}^{N}\left\{x\in B(x_{i},\ell_{N})\right\}\right)\leq N\mu(B(x,\ell_{N}))\asymp N\ell_{N}^{d}\to 0,\quad\text{ as }N\to\infty.$
@@ B.3 Kernel Ridgeless Regression with Benign Overfitting
Under Assumptions 1-4, and suppose either
@@
  • $k_{c_{N},\gamma_{N}}(x,x^{\prime})\coloneqq\tilde{k}(x,x^{\prime})+c_{N}\check{k}_{\gamma_{N}}(x,x^{\prime})$, where $\tilde{k}$ is a universal kernel, $\check{k}_{\gamma_{N}}$ is the Laplace kernel with bandwidth $\gamma_{N}>0$, $c_{N}\to 0$, $Nc_{N}^{4}\to\infty$, and $\gamma_{N}\leq N^{-3/d}(7\ln N)^{-1}$.
Then for any $\delta\in(0,1)$, there exist constants $C_{0},N_{0},\alpha>0$, for any $N\geq N_{0}$, define the uniform upper confidence bound as $U_{N}^{\delta}\coloneqq C_{0}\delta^{-1}N^{-\alpha}$, the following holds
@@
Therefore, for any threshold $\tau\in(0,0.98)$,
$\liminf_{N\to\infty}\mathbb{P}\left(\mathbb{E}_{D_{N}}\left[|f_{N}(x)|\geq\tau\right]\right)\geq 1-\delta,\quad\text{ for all }x\in\mathcal{C},$
which indicates that any hallucination detection criterion with a fixed $\tau$ fails in the correlation regions.
@@
$\mathcal{E}_{N}=O_{\mathbb{P}}\left(N^{-2\alpha}\right)\quad\text{ as }N,d\to\infty.$
@@
The proof is completed by combining the above result with Proposition 9 and Proposition 10. ∎
Appendix C Neural Kernels
Neural Tangent Kernels.
For an over-parametrized neural network of most architectures (e.g., multi-layer perceptron (MLP), residual network (ResNet), convolutional neural network (CNN), Transformer) under standard initialization (also known as the LeCun initialization) or neural tangent parametrization (jacot2018neural), the training dynamics of its output $f:\mathbb{R}^{d}\to\mathbb{R}$ can be tracked by the kernel gradient descent